# Translation from the Information Bottleneck Perspective: an Efficiency Analysis of Spatial Prepositions in Bitexts

### 20 Mar 2026 v1

@@ 5.2 Estimation of similarity between spatial relations evoked by prepositions
The correlation between judgements of word sense relatedness and computational representations relatedness has been widely investigated using both static embeddings (Hill et al., 2015), contextual embeddings and Large Language Models (Nair et al., 2020; Marjieh et al., 2022). Specifically, how context modulates the perceived meaning similarity between word pairs in written sentences has been graded by human subjects in many experiments (Huang et al., 2012; Pilehvar and Camacho-Collados, 2019; Armendariz et al., 2020; Vulić et al., 2020). However, most of these studies focus on nouns, adjectives or verbs. Our pilot study, conversely, was concerned with relations between entities, and the prepositions that mark them. Results presented in Table 1 indicate that a simple cosine similarity does not strongly correlate with our collected human sortings, whereas the low-rank projection model achieves a high Spearman correlation of $\rho=0.7800\pm 0.0622$. Interestingly, low-rank projection models achieve the highest correlation scores with $D=5$. This suggests that higher-dimensional projections likely lead to overfitting. These findings indicate that low-rank projections may be more effective than standard cosine similarity for predicting contextual similarities in spatial relations.
Aside from modelling, we acknowledge that our subset of 30 spatial prepositions in context does not guarantee cognitive representativeness of the domain of spatial relations. We also acknowledge that this subset is not fully representative of all linguistic means available for encoding spatial relations. While we focused on prepositions as a first step, alternative encoding mechanisms must be accounted for in future research, such as interactions between the preposition and the verb (see for example the verb-framing typologies from Talmy (1983) and Aurnague (2012)).
@@ 5.3 Efficiency of attested systems
In Figure 4 we observe that attested translations of spatial prepositions lie closer to optimal systems than random ones. To further investigate whether this is a feature of ”reasonable” translations, we generated counterfactual translations with varying degrees of perturbation. The results from Figure 5 indicate that as the degree of perturbation increases, the deviation to optimality increases. We believe that this comparison would benefit from including another baseline, such as a surjective mapping from meanings to terms, as it has been used in Strickland and Chemla (2018). This is equivalent to removing the non-deterministic assumption on the encoder $q$ (Strouse and Schwab, 2017). Future work should also test how altering the similarity geometry affects Accuracy. Nevertheless, these results suggest that attested systems exhibit a pressure for efficiency under the IB framework.
We conclude by addressing why attested translations deviate from optimality. This question lies at the crossroads of research on language efficiency and translation process (Teich et al., 2020). The translation process is indeed affected by the fluency of the translator (Lim et al., 2024), but especially stylistic incentives with respect to our case-study. This is illustrated by observations from the corpus: while German and Serbian translations of Le Tour du monde en 80 jours remain stylistically conservative, the English version acts as a ’free’ translation with added details. This divergence is clear in the original corpus, where LIFAT et al. (2024) found and explicitly annotated changes in meaning only in the English version. From an information-theoretic perspective, these changes in meaning represent a deviation from strict communicative efficiency, as the translator prioritised semantic elaboration or expressivity over the compact encoding predicted by the IB optimum.
@@ 6 Conclusion
This work proposes a novel methodology for analysing communicative efficiency via translation data. We performed an IB analysis on spatial prepositions across three translations of a French novel. Our results suggest that the translation of spatial prepositions exhibits pressure toward optimal trade-offs between informativity and complexity, consistent with the IB literature, and more broadly, recent evidence from language efficiency. Future work should extend this methodology to typologically distant language pairs to test whether this efficiency pressure is universal. We believe this is a plausible way to scale beyond fully elicited naming datasets while staying grounded in natural speech.
@@ 7 Ethics Statement
Written informed consent was obtained from all participants. Pile-sorting data consisted only of lists of sentences. These data were fully anonymised at the point of collection, such that no personal data was retained; the data from the pile-sorting experiment therefore fall outside the scope of the General Data Protection Regulation (GDPR, Regulation (EU) 2016/679). Given the minimal-risk nature of this feasibility study (involving no sensitive information and no identifiable data) institutional ethics committee approval was not sought.
@@ Appendix A Prompt for French to English few-shot alignment
2. Zero-alignment: if you cannot find an adequate spatial term in the target sentence corresponding to the translation of the source spatial term, don’t wrap anything in ¡loc¿ from the target sentence and return ‘None‘ for the normalized form.
3. If the meaning of the source spatial term is translated only by a verb or a noun in the target sentence, don’t wrap anything in ¡loc¿ from the target sentence and return ‘None‘ for the normalized form.
@@ Appendix C Pairwise similarity modelling
Appendix D Close-up of random encoders in the information plane
Figure D-1: Attested translations are displayed as small green stars, and 100 000 random encoders as small grey dots.
Appendix E Dimension reduction of the embedding space
Figure E-1: t-SNE visualisation of the 30 K-means centroids (computed from stacked final-layer representations of xlm-roberta-large) and embeddings of the most frequent French prepositions

### 05 Apr 2026 v2

@@ 5.2 Estimation of similarity between spatial relations evoked by prepositions
The correlation between judgements of word sense relatedness and computational representations relatedness has been widely investigated using both static embeddings (Hill et al., 2015), contextual embeddings and Large Language Models (Nair et al., 2020; Marjieh et al., 2022). Specifically, how context modulates the perceived meaning similarity between word pairs in written sentences has been graded by human subjects in many experiments (Huang et al., 2012; Pilehvar and Camacho-Collados, 2019; Armendariz et al., 2020; Vulić et al., 2020). However, most of these studies focus on nouns, adjectives or verbs. Our pilot study, conversely, was concerned with relations between entities, and the prepositions that mark them.
Results presented in Table 1 indicate that a simple cosine similarity does not strongly correlate with our collected human sortings, whereas the low-rank projection model achieves a high Spearman correlation of $\rho=0.7800\pm 0.0622$. Interestingly, low-rank projection models achieve the highest correlation scores with $D=5$. This suggests that higher-dimensional projections likely lead to overfitting. These findings indicate that low-rank projections may be more effective than standard cosine similarity for predicting contextual similarities in spatial relations.
Aside from modelling, we acknowledge that our subset of 30 spatial prepositions in context does not guarantee cognitive representativeness of the domain of spatial relations. However, results from Figure 3(a) and 3(b) suggest that colexification does not explain alone the similarity structure of spatial relations. We assume that modelling the semantic similarity of spatial relation in a continuous $D$-dimensional space accounts for the graded similarity space we observed, and that a sufficiently large bitext should provide broad coverage of the domain.
However, a lexical bias may arise during the pile sorting task, where shared word forms inflate human judgements of semantic similarity. As a result, similarity ratings may partly reflect surface-form overlap rather than purely relational semantics. This introduces a potential language-specific skew, whereby the lexical structure of French influences perceived relatedness. To mitigate this effect, future work should control for form overlap in the stimuli and compare across source languages, in order to quantify the contribution of lexical structure to similarity judgements.
We also acknowledge that this subset is not fully representative of all linguistic means available for encoding spatial relations. While we focused on prepositions as a first step, alternative encoding mechanisms must be accounted for in future research, such as interactions between the preposition and the verb (see for example the verb-framing typologies from Talmy (1983) and Aurnague (2012)).
@@ 5.3 Efficiency of attested systems
In Figure 4(a) we observe that attested translations of spatial prepositions lie closer to optimal systems than random ones. To further investigate whether this is a feature of “reasonable” translations, we generated counterfactual translations with varying degrees of perturbation. The results from Figure 4(b) indicate that as the degree of perturbation increases, the deviation to optimality increases. We believe that this comparison would benefit from including another baseline, such as a surjective mapping from meanings to terms, as used by Strickland and Chemla (2018). This is equivalent to removing the non-deterministic assumption on the encoder $q$ (Strouse and Schwab, 2017). Future work should also test how altering the similarity geometry affects Accuracy. Nevertheless, these results suggest that attested systems exhibit a pressure for efficiency under the IB framework.
We conclude by addressing why attested encoders deviate from optimality, and the confounds inherent to translation data. This question lies at the crossroads of research on language efficiency and translation process (Teich et al., 2020). Translating involves genre conventions and notably stylistic incentives. This is illustrated by observations from the corpus: while German and Serbian translations of Le Tour du monde en 80 jours remain stylistically conservative, the English version acts as a “free” translation with added details. This divergence is clear in the original corpus, where LIFAT et al. (2024) found and explicitly annotated changes in meaning, but were found only in the English version. From an information-theoretic perspective, these changes in meaning represent a deviation from strict communicative efficiency, as the translator prioritised semantic elaboration or expressivity over the compact encoding predicted by the IB optimum. Yet, a conservative translator might also imitate some of the lexical choices of the French original despite what would be natural in their language (Volansky et al., 2015; Lim et al., 2024). The alignment of preposition-to-preposition we performed (see section 4.1) is by design too coarse to capture the aforementioned phenomena. Therefore we argue that IB analysis of translation data must narrow the scope to relexification of terms of interest, such as prepositions, in order to capture only one aspect of the translation process.
@@ 6 Conclusion
This work proposes a novel methodology for analysing communicative efficiency via translation data. We performed an IB analysis on spatial prepositions across three translations of a French novel. Our preliminary results suggest that the translation of spatial prepositions exhibits pressure toward optimal trade-offs between informativity and complexity, consistent with the IB literature, and more broadly, recent evidence from language efficiency. An important next step is to apply this approach to typologically diverse, genealogically distant languages, allowing us to better separate general efficiency pressures from patterns arising through shared linguistic ancestry. We believe we presented a plausible way to extend to other semantic domains, other linguistic forms than prepositions, and scale beyond fully elicitepd naming datasets while staying grounded in natural speech.
@@ 7 Ethics Statement
Written informed consent was obtained from all participants. Pile-sorting data consisted only of lists of sentences. These data were fully anonymised at the point of collection, such that no personal data were retained; the data from the pile-sorting experiment therefore fall outside the scope of the General Data Protection Regulation (GDPR, Regulation (EU) 2016/679). Given the minimal-risk nature of this feasibility study (involving no sensitive information and no identifiable data) institutional ethics committee approval was not sought.
8 Acknowledgements
We sincerely thank our colleagues at Laboratoire d’InfoRmatique en Image et Systèmes d’information for participating in the pile-sorting experiment. We further thank the participants of the SIGTYP workshop for their stimulating discussions and valuable feedback during the presentation of this work. We are also grateful to the anonymous reviewers for their constructive suggestions, which helped us improve this paper.
@@ Appendix A Prompt for French to English few-shot alignment
2. Zero-alignment: if you cannot find an adequate spatial term in the target sentence corresponding to the translation of the source spatial term, don’t wrap anything in ¡loc¿ from the target sentence and return ”None” for the normalized form.
3. If the meaning of the source spatial term is translated only by a verb or a noun in the target sentence, don’t wrap anything in ¡loc¿ from the target sentence and return ”None” for the normalized form.
@@ Appendix C Pairwise similarity modelling
Appendix D Dimension reduction of the embedding space
Figure D-1: t-SNE visualisation of the 30 K-means centroids (computed from stacked final-layer representations of xlm-roberta-large) and embeddings of the most frequent French prepositions